# DPDFNet: Boosting DeepFilterNet2 via Dual-Path RNN

### v1

@@ IV-D Results
Turning to our proposed DPDFNet architectures, all variants consistently outperform the baseline models. Moreover, performance scales positively with the number of blocks (i.e., $k\in\{2,4,8\}$), with the largest model, DPDFNet-8, achieving the best overall results. When compared to strong causal baselines such as DPCRN, DPDFNet-8 delivers consistently higher fidelity and perceptual quality. Even against substantially larger waveform-domain models like DEMUCS and CleanUNet, DPDFNet variants achieve superior enhancement performance while requiring significantly fewer parameters and computational resources. Overall, the DPDFNet family compares favorably with both lightweight and heavyweight alternatives.
For completeness, although we consider the two benchmarks mentioned in Section III-B to be less representative of real-world conditions, we chose the DNS4 blind test set [dns4] because it provides longer, 10-second audio samples. We therefore evaluated our proposed models on DNS4 and compared their DNSMOS metrics with those of DeepFilterNet2/3. The results are shown in Table II.

### v2

@@ IV-D Results
Turning to our proposed DPDFNet architectures, all variants consistently outperform the baseline models. Moreover, performance scales positively with the number of blocks (i.e., $k\in\{2,4,8\}$), with the largest model, DPDFNet-8, achieving the best overall results. When compared to strong causal baselines such as DPCRN, or even to the latest and most advanced models like FastEnhancer-L, DPDFNet-8 consistently delivers higher fidelity and perceptual quality. Even against substantially larger waveform-domain models like DEMUCS and CleanUNet, DPDFNet variants achieve superior enhancement performance while requiring significantly fewer parameters and computational resources. Overall, the DPDFNet family performs favorably compared to both lightweight and heavyweight alternatives.
For completeness, although we consider the two benchmarks mentioned in Section III-B to be less representative of real-world conditions, we chose the DNS4 blind test set [5] because it provides longer, 10-second audio samples. We therefore evaluated our proposed models on DNS4 and compared their DNSMOS metrics with those of DeepFilterNet2/3. The results are shown in Table II.